OpenStudy (unklerhaukus):

is \[(-1)^{1-n}=(-1)^{n+1}\]

OpenStudy (unklerhaukus):

\[=(-1)^{n-1}?\]

myininaya (myininaya):

which equality are you wanting? are there any restrictions on n?

OpenStudy (unklerhaukus):

sorry n is intergers;

OpenStudy (anonymous):

Plug in values for n and see... If n =0 you get what? -1 for both?

OpenStudy (anonymous):

Just plug in n values.

OpenStudy (anonymous):

Let 1 - n be a \[(-1)^a = (-1)^{a+2n}\] If a is odd, then a + 2n is odd since odd + even = odd. If a is even, then a + 2n is even since even + even = even.

myininaya (myininaya):

i'm still looking at \[(-1)^{1-n}=(-1)^{n+1}\]

OpenStudy (unklerhaukus):

so for n={1,2,3,...} \[(-1)^{n+1} = (-1)^{n-1} = (-1)^{1-n} = (-1)^{1+n}\] ?

myininaya (myininaya):

lol i got confused

myininaya (myininaya):

just plug in values like male said to see what happens with whatever you are checking is true

OpenStudy (anonymous):

If n is odd, then n + 1 is even n -1 is also even 1 - n is even since odd - odd = even

myininaya (myininaya):

money posted some good stuff too

OpenStudy (unklerhaukus):

well i plugged in values and that told me the expressions are all equivalent , i was just making sure that i can substitute these terms at will.

myininaya (myininaya):

did you look at what money wrote?

OpenStudy (unklerhaukus):

does anything change if if the \[(-1)^{n+1}\]is in a summation over n \[ \sum_{n=1}^∞ \]

myininaya (myininaya):

\[\text{ if n is odd then n+1 and n-1 and 1-n are even}\] \[(-1)^{n+1}=(-1)^{n-1}=(-1)^{1-n}\] all these powers are even integers which means we have 1=1=1 so what if n is even then n+1 and n-1 and 1-n are odd \[(-1)^{n-1}=(-1)^{1-n}=(-1)^{n+1}\] since -1=-1=-1

OpenStudy (unklerhaukus):

im guessing it doesnt change anything

OpenStudy (unklerhaukus):

thank you all for your patients